malcolm paid 28$ per tie for a recent purchase of three ties the price of one tie was twice as much as another and the remaining tie cost 24$ whjat were the prices of the other two ties
Answer by oberobic(2304)

3 ties cost $28. (Note the '$' is placed in front of the value.)
.
3 ties could be defined: a, b, c.
.
one tie was twice as much as one of the others
b = 2a
.
the remaining tie cost $24
c = 24
.
a +2a + 24 = 28
3a = 4
a = 4/3
a = $1.33
.
b = 2a
b = 8/3
b = $2.67
.
Check your work.
a + b + c = ?
24 + 1.33 + 2.67 = 24 + 4 = 28
Correct.
.
The two other ties costs $1.33 and $2.67 each.
.
Notice that I rounded 4/3 to 1.33 and, to make the total = $4, I rounded 8/3 to 2.67.
.
Done.